OpenStudy (31356):

A sphere has a surface area equal to 125 square feet. Find the length of the diameter of the sphere. Round your answer to the nearest hundredth. Use 3.14 to approximate pi. MEDAL REWARDED!!

OpenStudy (31356):

A) 3.15 B) 9.91 C)6.31 D)19.81

OpenStudy (31356):

@jdoe0001

OpenStudy (31356):

@tomhue

OpenStudy (jdoe0001):

\(\bf \textit{surface area of a sphere}=4\pi r^2\qquad r=radius=\frac{diameter}{2}\\ \quad \\ 125=4\pi r^2\) solve for "r"

OpenStudy (anonymous):

I think its B) but i might be wrong so check it